Venz Company manufactures LED TVs, stereos and speakers, using a common parts inventory of chassis, LCD...

  1. Venz Company manufactures LED TVs, stereos and speakers, using a common parts inventory of chassis, LCD Screens, speakers, power supplies and various internal electronics. Parts are in limited supply and you must determine the most profitable mix of products to build. The table below contains the part requirement per product, part inventory available, and contribution margin per product.

Part Name

LED TV

Stereo

Speaker

Available

Chassis

1

1

0

450

LED Screen

1

0

0

250

Speaker

2

2

0

800

Power Supply

1

1

0

450

Internal Electronics

2

1

1

600

Contribution Margin

$75

$50

$35

Required:

What is the most profitable product mix?
Answer #1

Maximum Contribution Margin is in LED TV.

Hence maximum critical resources should be allocated to LED TV

Part Name Inventory Quantity Required per LED TV Maximum LED TV that can be produced
Chasis 450 1 450
LED Screen 250 1 250
Speaker 800 2 400
Power Supply 450 1 450
Internal Electronics 600 2 300

Hence , 250 LED TVs should be produced

Balance Inventory:

Chasis:450-250=200

LED Screen=250-250=0

Speaker=800-(2*250)=300

Power Supply:450-250=200

Internal Electronics:600-(2*250)=100

Next highest contribution margin is provided by Stereo

Part Name Inventory Quantity Required per Stereo Maximum Number Stereo that can be produced Number of parts used for 100 Stereos Balance Inventory
Chasis 200 1 200 100 100
LED Screen 0 0 0 0
Speaker 300 2 150 200 100
Power Supply 200 1 200 100 100
Internal Electronics 100 1 100 100 0

Number of Stereos Produced =100

Since Internal Electronics Inventory =0, Speakers cannot be produced.

Most Profitable Mix:

Product Quantity Unit Contribution Total Contribution
LED TV 250 $75 $18,750
STEREO 100 $50 $5,000

Total Contribution: 18750+5000=$23,750